# Evaluating Mathematical Reasoning Beyond Accuracy

### v1

@@ 5.1 Evaluating Reasoning Quality of LLMs Specialized in Math
The models selected for measurement are two mainstream LLMs specialized in math: Abel (Chern et al., 2023) and WizardMath (Luo et al., 2023), with different scales. We also report the results of LLaMA-2 (Touvron et al., 2023) naive fine-tuned on PRM800K (approximately 6K solutions that reach the correct final answers) for comparison. We evaluate the performance of the LLMs on MATH and sample 100 solutions for each problem to reduce evaluation noise and sampling randomness. Specifically, the sampling temperature is set to 0.6, and the top-p value is set to 0.95. The maximum sample length is set to 2048 tokens. All solutions are scored by ReasonEval.
Our analysis includes two aspects: (1) False positive rate: it refers to the proportion of solutions that have the correct final answers but contain incorrect steps among all solutions with correct final answers. (2) The redundancy of solutions. We also analyze the first error location in reasoning steps and present the results and analysis in Appendix B.1.
5.1.1 False Positive Rate
| Model | Base Model | Methods | Acc. (%) | FPR (%) |
| LLaMA2-13B-PRM800K | LLaMA2 | SFT | 7.4 | 40.6 |
| LLaMA2-70B-PRM800K | 14.7 | 21.8 | ||
| Abel7B-001 | LLaMA2 | SFT | 12.4 | 31.8 |
@@ 5.1.1 False Positive Rate
| Abel70B | 25.7 | 20.0 | ||
| Abel7B-002 | Mistral | 30.0 | 22.8 | |
| WizardMath7B-V1.0 | LLaMA2 | SFT+RLHF | 8.9 | 34.4 |
@@
| WizardMath70B | 18.7 | 16.5 | ||
| WizardMath7B-V1.1 | Mistral | 31.0 | 16.7 |
Figure 4: Correlation between the final-answer accuracy and the false positive rate.
Figure 5: Box-and-whisker plots of (upper) and (lower). The boundaries of the whiskers are based on the 1.5 interquartile range.
First, we compare the validity scores of reasoning steps from different models in Figure 5. For solutions with the correct answers but wrong steps, the validity scores should behave similarly to solutions with the incorrect answers. So we can use it to estimate the threshold and calculate the false positive rate. A rough estimation for the threshold is between 0.15 and 0.35. To further ensure a relatively accurate estimation, we check the false positive rate of Abel13B and WizardMath13B by humans, sampling one solution for each problem. Combining above information, the threshold is set to 0.25. We present the main results in Table 5 and Figure 4.
@@
The model size and base model affect the false positive rate significantly. When comparing LLaMA13B-PRM800K with LLaMA70B-PRM800K, although using the same training data, the distinction in false positive rates is large (40.6% vs. 21.8%). It indicates the importance of the model scale. The base model of Mistral (Jiang et al., 2023) also achieves a lower false positive rate than LLaMA2 (22.8% vs. 31.8%, 16.7% vs. 34.4%). Overall, the scaling law still applies to this field.
@@ 5.2 Improving Reasoning Quality by Selecting Training Data
| Filter | #Training | Acc. | Val. | Red. | #Token |
@@
Setup
The MMIQC (Liu & Yao, 2024) dataset consists of a mixture of processed web data and synthetic question-response pairs used to enhance the reasoning capabilities of LLMs. We randomly sample 10K unique responses generated by GPT-3.5-turbo on MATH from the dataset. We then filter it using , specifically removing samples with validity scores below 0.5 or redundancy scores above 0.15. We also combine these two conditions. We SFT the mistral-7b in different subsets with 3 epochs and observe the performance on MATH. To compare, we randomly sample the same amount of training data three times and report the average performance.
Results
@@ 6 Related Work
The technique used to assess reasoning steps automatically can be broadly divided into three groups: (1) Embedding-based methods: Golovneva et al. (2022) propose several metrics and uses the embedding-based calculation among hypothesis steps, reference steps, and problems to represent them; (2) Parsing-based methods: this approach aims to parse steps into structured forms, like ‘subject-verb-object’ frames (Prasad et al., 2023) or symbolic proofs (Saparov & He, 2022). However, it presents challenges for complex datasets such as MATH due to the intricate logic involved; (3) Prompting-based methods: due to the generalization of LLMs, given tailored prompts to SOTA LLMs, they can check the solutions and find the potential errors (Tyen et al., 2023; Zeng et al., 2024).
@@ 7 Conclusion
In this work, we develop ReasonEval, a new metric to evaluate the quality of reasoning steps from the perspective of correctness and efficiency. Our experiments show that ReasonEval can effectively detect different types of errors and achieves competitive performance compared to other methods. With ReasonEval, we find the inconsistency between final-answer accuracy and the quality of reasoning steps. Additionally, we discover its efficacy in selecting training data.
Appendix A Supplements for Meta-Evaluations
A.1 Examples of Solutions Generated by Different LLMs
We present examples of solutions generated by different LLMs in Figure 9, including LLaMA-PRM800K, Abel and WizardMath.
A.2 Training Details for ReasonEval
The ReasonEval model is trained on NVIDIA A100 GPUs (4 GPUs for the 7B models, 16 GPUs for the 34B models) using the supervised fine-tuning (SFT) framework provided by Dai et al. (2023). Specifically, the model is trained in 1 epoch. We set the batch size to 64 and maximum sequence length to 2048. We use a peak learning rate 1e-6 with 8 warmup steps and cosine learning rate decay to 0. We use AdamW (Loshchilov & Hutter, 2018) as our optimizer with and weight decay of 0.1. Only the last token of a step is taken into account in the loss function, as it contains the full information about the step.
A.3 Settings for Evaluators
@@
A.4 Addition Information for Meta-Evaluations by Perturbation
In Figure 3, for swapping solutions, the redundancy score of the original solution is higher than others. This observation may stem from the fact that the first step (i.e., the observed step) typically involves analyzing the problem from scratch, making it more likely to involve restating the problems, which may be considered an unproductive or unnecessary step for the overall solution.
We present the results of solution-level scores in Figure 8. For solution-level scores, the tendencies discussed in §4.2 remain unchanged, but the distinction between the hypothesis and reference decreases, as the solution-level score can be disturbed by other steps. For reference-free evaluation, it adds difficulty in finding the specific step.
subfigure
subfigure
Appendix B Supplements for Evaluating LLMs
B.1 Location of the First Error Step
We examine the characteristics of the first error location identified by the validity score. We set the threshold to 0.25 for high recall. We present the main results in Figure 8. While there is significant noise in them, we can still observe general trends. For reasoning steps with incorrect results, the model is more prone to making mistakes at an early stage, leading to false results, especially when the lengths of reasoning steps are longer. For reasoning steps with correct results, the distribution is much flatter, indicating these errors hidden in reasoning steps can happen at any location.

### v2

@@ 5.1 Evaluating Reasoning Quality of LLMs Specialized in Math
The models selected for measurement are two mainstream LLMs specialized in math: Abel (Chern et al. 2023) and WizardMath (Luo et al. 2023), with different scales. We also report the results of LLaMA-2 (Touvron et al. 2023) naive fine-tuned on PRM800K (approximately 6K solutions that reach the correct final answers) for comparison. We evaluate the performance of the LLMs on MATH and sample 100 solutions for each problem to reduce evaluation noise and sampling randomness. Specifically, the sampling temperature is set to 0.6, and the top-p value is set to 0.95. The maximum sample length is set to 2048 tokens. All solutions are scored by ReasonEval.
Our analysis includes two aspects: (1) False positive rate: it refers to the proportion of solutions that have the correct final answers but contain incorrect steps among all solutions with correct final answers. (2) The redundancy of solutions.
False Positive Rate
| Model | Acc. (%) | FPR (%) |
| LLaMA2-13B-PRM800K | 7.4 | 40.6 |
| LLaMA2-70B-PRM800K | 14.7 | 21.8 |
| Abel7B-001 | 12.4 | 31.8 |
@@
| Abel70B | 25.7 | 20.0 |
| Abel7B-002 | 30.0 | 22.8 |
| WizardMath7B-V1.0 | 8.9 | 34.4 |
@@
| WizardMath70B | 18.7 | 16.5 |
| WizardMath7B-V1.1 | 31.0 | 16.7 |
We set the threshold to 0.25 for the validity scores and calculate the false positive rate. We describe the details of choosing the threshold value in Appendix E. We present the main results in Table 4. We also include the results in Figure 4 with the information of the tested models. We highlight three key takeaways:
@@
The model size and base model affect the false positive rate significantly. When comparing LLaMA13B-PRM800K with LLaMA70B-PRM800K, although using the same training data, the distinction in false positive rates is large (40.6% vs. 21.8%). It indicates the importance of the model scale. The base model of Mistral (Jiang et al. 2023) also achieves a lower false positive rate than LLaMA2 (22.8% vs. 31.8%, 16.7% vs. 34.4%). Overall, the scaling law still applies to this field.
@@ 5.2 Improving Reasoning Quality by Selecting Training Data
Setup
The MMIQC (Liu and Yao 2024) dataset consists of a mixture of processed web data and synthetic question-response pairs used to enhance the reasoning capabilities of LLMs. We randomly sample 10K unique responses generated by GPT-3.5-turbo on MATH from the dataset. We then filter it using , specifically removing samples with validity scores below 0.5 or redundancy scores above 0.15. We also combine these two conditions. We SFT the mistral-7b in different subsets with 3 epochs and observe the performance on MATH. To compare, we randomly sample the same amount of training data three times and report the average performance.
Results
| Filter | #D | Acc. | Val. | Red. | #Token |
@@ 6 Related Work
The technique used to assess reasoning steps automatically can be broadly divided into three groups: (1) Embedding-based methods: Golovneva et al. (2022) propose several metrics and uses the embedding-based calculation among hypothesis steps, reference steps, and problems to represent them; (2) Parsing-based methods: this approach aims to parse steps into structured forms, like ‘subject-verb-object’ frames (Prasad et al. 2023) or symbolic proofs (Saparov and He 2022). However, it presents challenges for complex datasets such as MATH due to the intricate logic involved; (3) Prompting-based methods: due to the generalization of LLMs, given tailored prompts to SOTA LLMs, they can check the solutions and find the potential errors (Tyen et al. 2023; Zeng et al. 2024).
@@ 7 Conclusion
In this work, we develop ReasonEval, a new metric to evaluate the quality of reasoning steps from the perspectives of correctness and efficiency. We construct a meta-evaluation testbed and show ReasonEval using the optimal design can consistently outperform baseline methods by a large margin. Additionally, we empirically demonstrate the strong generalization ability of ReasonEval. With ReasonEval, we find an inconsistency between final-answer accuracy and the quality of reasoning steps. We also show its efficacy in selecting training data.
Appendix A Ablation Studies on the Scoring Scheme
| Solution-level | Step-level | |||
| -AUC | -AUC | -AUC | -AUC | |
| ReasonEval | 79.5 | 74.0 | 80.0 | 67.1 |
| ReasonEval | 81.9 | 80.3 | 83.9 | 77.5 |
| ReasonEval | 85.1 | 76.5 | 85.7 | 73.5 |
| ReasonEval | 84.3 | 82.9 | 90.5 | 85.9 |
| ReasonEval | 86.2 | 83.5 | 90.5 | 81.3 |
| ReasonEval | 87.5 | 84.3 | 89.5 | 83.1 |
| ReasonEval | 90.8 | 89.7 | 92.8 | 89.6 |
| MR-MATH-invalid | MR-MATH-redundant | |||||
| MIN | AM | GM | MAX | AM | GM | |
| ReasonEval | 80.0 | 74.0 | 73.5 | 62.8 | 65.5 | 66.6 |
| ReasonEval | 83.9 | 80.3 | 78.3 | 65.6 | 66.2 | 66.2 |
| ReasonEval | 85.7 | 76.5 | 80.2 | 63.4 | 64.4 | 64.5 |
| ReasonEval | 90.5 | 82.9 | 82.5 | 63.0 | 63.5 | 63.2 |
| ReasonEval | 90.5 | 83.5 | 85.6 | 63.6 | 64.7 | 65.3 |
| ReasonEval | 89.5 | 84.3 | 87.9 | 64.8 | 65.4 | 65.4 |
| ReasonEval | 92.8 | 89.7 | 90.3 | 62.7 | 64.2 | 64.2 |
Validity Scores
Aggregation Methods
We explore three different methods for aggregating step-level scores: min/max, the arithmetic mean (AM), and the geometric mean (GM). We compare the performance in Table 7. For the validity scores, the min operation consistently outperforms both the AM and GM by a large margin. Conversely, for the redundancy scores, the AM and GM slightly outperform the max operation. This discrepancy may be due to the fact that the redundancy score estimations are less accurate compared to validity scores for ReasonEval, and averaging operations help to reduce noise. Given the trade-offs and considering the potential future advancements in model capabilities, we ultimately choose the min/max operation, as also taken by Golovneva et al. (2022) and Prasad et al. (2023).
Appendix B Details for ReasonEval
Training Details
The ReasonEval model is trained on NVIDIA A100 GPUs (4 GPUs for the 7B models, 16 GPUs for the 34B models) using the supervised fine-tuning (SFT) framework provided by Dai et al. (2023). Specifically, the model is trained in 1 epoch. We set the batch size to 64 and maximum sequence length to 2048. We use a peak learning rate 1e-6 with 8 warmup steps and cosine learning rate decay to 0. We use AdamW (Loshchilov and Hutter 2019) as our optimizer with and weight decay of 0.1.
Splitting Solutions
[…]
Appendix C Examples of Solutions Generated by Different LLMs
We present examples of solutions generated by different LLMs in Figure 5, including LLaMA-PRM800K555It represents the LLaMA-2 naive fine-tuned on PRM800K (approximately 6K solutions that reach the correct final answers), reflecting the language style of PRM800K., Abel and WizardMath.
Appendix D Settings for Evaluators
@@
Appendix E Threshold Selection for False Positive Rate Estimation
First, we compare the validity scores of reasoning steps from different models in Figure 3. For solutions with the correct answers but wrong steps, the validity scores should behave similarly to solutions with the incorrect answers. So we can use it to estimate the threshold and calculate the false positive rate. A rough estimation for the threshold is between 0.15 and 0.35. To further ensure a relatively accurate estimation, we check the false positive rate of Abel13B and WizardMath13B by humans, sampling one solution for each problem. Combining above information, the threshold is set to 0.25.